is the ordered pair a solution to the equation 4x+2=24; (4,4) show all work to find the answer
Answer by FrankM(1040)

4x+2=24
4x=22
x=5.5
the pair is not a solution.
alternate answer -
4(4)+2=24
16+2=24
18=24
no. the 4 we entered into the equation produces an invalid result.